# From Handcrafted Features to LLMs: A Brief Survey for Machine Translation Quality Estimation

### v1

@@ III-B1 Classic Deep Learning Methods
With the advancement of QE, the advent of word embeddings [turian2010word, mikolov2013linguistic, mikolov2013efficient, pennington2014glove] and neural machine translation (NMT) [bahdanau2014neural, sutskever2014sequence] technologies has led some researchers to apply neural networks to QE tasks. The progression from initially utilizing neural networks for feature extraction to the emergence of fully neural network-based QE systems has greatly enhanced the performance of QE systems.
@@
Besides utilizing the handcrafted features from QuEst, SHAH et al. [shah2015shef, shah2015investigating] also employed additional word-level QE features extracted from Word2Vec [mikolov2013linguistic, mikolov2013efficient] embeddings and the similarity in the embedding space between source and target language words. They combined language model probabilities generated from trained continuous space models with these handcrafted features for sentence-level QE. Furthermore, Scarton et al. proposed word embedding features [scarton2016word], discourse features, and features extracted from pseudo-reference translations [scarton2015searching] for document-level QE. Inspired by their work, Chen et al. [chen2017improving] proposed the use of sentence embedding features and cross-entropy features to enhance the correlation of QE with human evaluations and investigated several factors affecting the performance of QE systems.
@@
Subsequently, some researchers explored the use of neural networks solely for feature extraction and QE. QUETCH [kreutzer-etal-2015-quality] is an early example of this approach, utilizing pretrained word representations and a Deep Neural Network (DNN) architecture. The QUETCH comprises an input layer, lookup tables, a multilayer perceptron (MLP), and an output layer. It feeds bilingual context representations through a fixed-size word window into the MLP, ultimately completing the word-level QE task via the output layer. However, its effectiveness did not match that of QUETCH+, which integrated additional baseline features. Expanding on QUETCH, Martins et al. [martins-etal-2016-unbabels] introduced a 200-unit bi-directional Gated Recurrent Unit (BiGRU) network and stacked feed-forward neural networks, subsequently incorporating POS tags for both source and target words to achieve the best performance of that time. Similar to QUETCH, Patel et al. [patel2016translation] made a switch from DNN to RNN, utilizing LSTM and GRU for extracting representations of bilingual sequences, and introduced sub-labels to address the challenge of label imbalance.
@@
While the QUETCH [kreutzer-etal-2015-quality] approach relies entirely on neural networks for feature extraction, it requires bilingual alignment information, which is typically obtained via statistical methods and is prone to significant errors. With the advancement of deep learning technologies, the trend in QE research has been gradually shifting towards completely neural network-based methods.
@@
In 2016, Kim et al. [kim-lee-2016-recurrent, kim2016recurrent] made the first attempt to use an NMT model for QE, proposing the inaugural purely neural approach for sentence-level, word-level, and phrase-level QE that does not require manually extracted features. In 2017, Kim et al. [kim2017predictor] presented a more in-depth study and named it the Predictor-Estimator (PredEst) model, which is a method to address the issues of expensive QE annotations and limited annotated QE data. It consists of two components: the predictor and the estimator. The predictor is a neural word prediction model trained using parallel corpora. It masks the target word, feeds the source language and the corrupted target language into bidirectional RNN (Bi-RNN), and predicts the probability distribution of the masked word. On the other hand, the estimator is a neural QE model trained on QE data, extracting QE feature vectors (QEFVs) and training them on a feedforward network. QEFVs are processed by FNN, RNN, or Bi-RNN to obtain hidden representations, which are then used to predict quality labels for sentence, phrase, or word-level tasks. Later, to train the model effectively, Kim et al. [kim-etal-2017-predictor] introduced stack propagation and multi-level task algorithms to improve the original method. In 2018, Ive et al. [ive2018deepquest] proposed the deepQuest framework for sentence-level and document-level QE, marking the first purely neural approach for document-level QE, which for the first time attempted to experiment with the outputs of both SMT and NMT. Upon testing, the framework proved to be faster and more cost-effective, and greatly improved the performance of the document-level QE framework.
@@
Martins et al. [martins2017pushing, martins2017unbabel] introduced a STACKEDQE system that stacked both linear and neural systems in the WMT17 word-level QE task, followed by combining APE with word-level QE to create an APEQE system. Ultimately, they merged these two systems to form the FULLSTACKEDQE system tailored for word-level QE and extended the FULLSTACKEDQE to sentence-level QE. These systems all achieved commendable results. Building on the approach of Martins et al. [martins2017pushing, martins2017unbabel], Hokamp et al. [hokamp2017ensembling] incorporated features that had been demonstrated to be effective for word-level QE into the input of an NMT system, thus proposing the APE-QE model. This unified APE with word-level QE models and achieved the best performance of the time in both APE and QE tasks.
@@
As the Transformer [DBLP:journals/corr/VaswaniSPUJGKP17] model has garnered significant success in the field of MT, Fan, Wang, et al. [DBLP:journals/corr/abs-1807-09433, wang2018alibaba] developed a Bilingual Expert model based on a bidirectional Transformer and a PredEst architecture that includes a word prediction module and a QE module. The word prediction module leverages prior knowledge obtained from pretraining on a large parallel corpus and the joint latent representations between the source language and the translation for token prediction, extracting a set of features. Then, they introduce mismatch features that measure the discrepancy between the prior knowledge obtained from well-trained Bilingual Expert and the targets in the QE dataset to train the QE module, which uses a bidirectional LSTM model, achieving SOTA performance at that time. Wang et al. [wang-etal-2020-hw-tscs] employed a pre-trained Transformer as the predictor and integrated Bottleneck Adapter Layers (BAL) for efficient transfer learning, with specific classifier and regressor as the estimator. They also conducted joint training for word and sentence-level tasks using a unified model and proposed a pseudo PE assisted QE method. This demonstrated the effectiveness of using pre-trained NMT models for transfer learning in QE tasks.
@@
However, Cui et al. [cui2021directqe] argued that the gap between data quality and training objectives in the PredEst framework hindered its ability to benefit from parallel corpora. Consequently, they proposed a framework called DirectQE, which includes a generator for creating pseudo QE data and a detector pre-trained with these pseudo data. This framework allows for pre-training using large parallel corpora and fine-tuning on real QE data, thereby addressing the issues inherent in the PredEst framework.
@@ III-B2 Incorporating Pre-trained Language Models Methods
With the emergence and development of pre-trained LMs such as ELMo [peters2018deep], BERT [devlin2018bert], XLM [lample2019cross], and XLM-R [conneau2019unsupervised, DBLP:journals/corr/abs-1911-02116], some studies have begun to integrate pre-trained LMs into QE models. This integration has enabled better extraction of quality vectors from source texts and translated texts, thereby enhancing the performance of QE systems.
@@
Kepler et al. [kepler2019unbabel] expanded OpenKiwi [kepler2019openkiwi] into a Transformer-based PredEst model, replacing the predictor with the pre-trained LMs BERT and XLM, and proposed an ensemble method using the POWELL technique to combine word-level and sentence-level predictions. Moreover, they suggested a simple technique for converting word labels into document-level predictions. Wu et al. [wu2020tencent] extended the OpenKiwi by integrating PredEst models based on XLM and Transformer in their submission to WMT20. The former predictor generates both masked and nonmasked representations, while the latter produces only nonmasked representations. The estimator is trained using either LSTM or Transformer, employing top-K and multi-head attention strategies to enhance sentence feature representation. Ranasinghe et al. [DBLP:journals/corr/abs-2011-01536] proposed TransQuest, a PredEst model aimed at reducing the dependence of sentence-level QE on large-scale parallel corpora. TransQuest does not use parallel data to pre-train the predictor but instead directly employs SOTA cross-lingual embedding models like XLM-R [conneau2019unsupervised, DBLP:journals/corr/abs-1911-02116] to encode the source and target sentences. It consists of two neural networks: MonoTransQuest (MTransQuest) and SiameseTransQuest (STransQuest). MTransQuest uses a single XLM-R model to encode concatenated source and target sentences, whereas STransQuest adopts a Siamese architecture, using separate XLM-R models for the source and translation. Both models use mean squared error loss as the objective function and have shown improved results with specific pooling strategies. Zerva et al. [zerva2021unbabel] used a pre-trained multilingual encoder combined with adapters, training multilingual models on the OpenKiwi [kepler2019openkiwi] PredEst and found that adapter tuning can resist overfitting. Additionally, they demonstrated that integrating uncertainty information and using out-of-domain data for pre-training can improve QE system performance.
@@
Zhou et al. [zhou2019source] primarily investigated the application of pre-trained translation models in QE and compared the effectiveness of a bilingual expert, ELMo, and BERT on QE tasks. Yankovskaya et al. [yankovskaya2019quality] contrasted two approaches: one using only BERT and LASER [artetxe2019massively] embeddings as features, and the other additionally incorporating log probability features from MT systems. Their research demonstrated the importance of the log probabilities from MT systems.
@@
In 2020, Rei et al. [rei2020comet] introduced COMET, a neural framework for training multilingual and adaptable MT evaluation models, which is often used for reference-based evaluation to generate predictive estimates of human judgments such as HTER, DA, and MQM. The COMET framework supports two different architectures: an Estimator model and a Translation Ranking model, both consisting of a cross-lingual encoder and pooling layers, with the fundamental difference being the training objective. The Estimator model, which is the most commonly used, is trained to directly regress to a quality score, while the Translation Ranking model is trained to minimize the distance between a “better” hypothesis and its corresponding reference translation and source language.
@@
In 2022, Rei et al. [rei2022cometkiwi] combined the strengths of COMET and OpenKiwi [kepler2019openkiwi] by connecting COMET with the PredEst architecture of OpenKiwi, and equipped it with a word-level sequence tagger and explanation extractor, forming COMETKIWI for QE. COMETKIWI pre-trained the model on metrics data with the learning objective proposed by UniTE model [wan2022unite], which incorporates reference translation into the training, serving as a form of data augmentation. Additionally, COMETKIWI proposed an interpretability method using attention and gradient information and further refined the influence of attention heads on predictions through a Head Mix module. COMETKIWI also demonstrated the effectiveness of few-shot learning, achieving significant improvements in model performance with only 500 samples.
@@ III-C1 Direct prediction based on content generated by LLMs
Kocmi and Federmann [kocmi2023large] proposed GEMBA, a GPT-based [brown2020language] translation quality assessment metric that uses single-step prompting and can be applied to scenarios with reference translations as well as to QE. They evaluated 9 different GPT models and concluded that only GPT-3.5 and larger models are capable of performing QE. GEMBA focuses on zero-shot prompting, with the authors employing 4 different prompt templates to execute quality assessments for both reference-based and nonreference-based translation modes. GEMBA predicts scores directly based on content generated by LLMs, assessing each segment independently, then averaging the scores of all segments to obtain a final system-level score, achieving SOTA performance at the system level, but lacking in segment-level analysis. To enhance the performance of LLMs in quality assessment, Lu et al. [lu2023error] introduced Error Analysis Prompting (EAPrompt), a new prompting method that combines Chain-of-Thought (CoT) [wei2022chain] with EA [lu2022toward]. Using ChatGPT, this approach predicts the degree and number of errors and provides a score based on the severity of those errors, creating MQM-like assessments. It achieved better results on CPT-3.5-turbo than GEMBA. Yang et al. [yang2023knowledge] introduced the Knowledge-Prompted Estimator (KPE), which is a CoT approach that integrates three single-step prompting techniques, using LLMs to predict fluency, word-level similarity [zhang2023implicit], and sentence-level similarity [yang2023teachersim], resulting in better performance for segment-level QE. Furthermore, KPE also demonstrated its advantages in terms of interpretability [tao2022crossqe].
@@ III-C2 Based on the generative probabilities of LLMs
Huang et al. [huang2023towards] leveraged various prompts and examples to obtain multiple generation probabilities for a certain source sentence and its corresponding translated sentence in the GPT-3.5 interface. They then calculated a more accurate measure of uncertainty as an assessment of the quality of the translated sentence by computing the mean and variance of these probabilities.
@@ III-C3 Leveraging LLMs to generate pseudo data
Xu et al. [xu2023instructscore] introduced INSTRUCTSCORE, a method that learns interpretable text generation metrics without the need for human-annotated scores, by constructing MQM-like data using the knowledge provided by GPT-4 [openai2023gpt4] to train the LLaMA model [touvron2023llama]. Moreover, Huang et al. [HUANG2024102022] also leveraged LLMs to corrupt reference sentences, subsequently generating a fluent sentence from the corrupted one and outputting it to obtain a noisy negative view. Since this approach does not require data annotation, it has strong generalization capabilities.
@@ III-C4 LLMs as the foundation for QE models
Gladkoff et al. [gladkoff2023predicting] fine-tuned LLMs using the OpenAI API interface to assess whether translations require editing. Similarly, Xu et al. [xu2023instructscore], as mentioned above, also utilized pseudo data generated by GPT-4 [openai2023gpt4] to train the LLaMA model [touvron2023llama].

### v2

@@ III-B1 Classic Deep Learning Methods
With the advancement of QE, the advent of word embeddings [20, 21] and neural machine translation (NMT) [22, 23] technologies has led some researchers to apply neural networks to QE tasks. The progression from initially utilizing neural networks for feature extraction to the emergence of fully neural network-based QE systems has greatly enhanced the performance of QE systems.
@@
Besides utilizing the handcrafted features from QuEst, SHAH et al. [24, 25] also employed additional word-level QE features extracted from Word2Vec [20] embeddings and the similarity in the embedding space between source and target language words. They combined language model probabilities generated from trained continuous space models with these handcrafted features for sentence-level QE. Furthermore, Scarton et al. proposed word embedding features [26], discourse features, and features extracted from pseudo-reference translations [27] for document-level QE. Inspired by their work, Chen et al. [28] proposed the use of sentence embedding features and cross-entropy features to enhance the correlation of QE with human evaluations and investigated several factors affecting the performance of QE systems.
@@
Subsequently, some researchers explored the use of neural networks solely for feature extraction and QE. QUETCH [29] is an early example of this approach, utilizing pretrained word representations and a Deep Neural Network (DNN) architecture. The QUETCH comprises an input layer, lookup tables, a multilayer perceptron (MLP), and an output layer. It feeds bilingual context representations through a fixed-size word window into the MLP, ultimately completing the word-level QE task via the output layer. However, its effectiveness did not match that of QUETCH+, which integrated additional baseline features. Expanding on QUETCH, Martins et al. [30] introduced a 200-unit bi-directional Gated Recurrent Unit (BiGRU) network and stacked feed-forward neural networks, subsequently incorporating POS tags for both source and target words to achieve the best performance of that time. Similar to QUETCH, Patel et al. [31] made a switch from DNN to RNN, utilizing LSTM and GRU for extracting representations of bilingual sequences, and introduced sub-labels to address the challenge of label imbalance.
@@
While the QUETCH [29] approach relies entirely on neural networks for feature extraction, it requires bilingual alignment information, which is typically obtained via statistical methods and is prone to significant errors. With the advancement of deep learning technologies, the trend in QE research has been gradually shifting towards completely neural network-based methods.
@@
In 2016, Kim et al. [32, 33] made the first attempt to use an NMT model for QE, proposing the inaugural purely neural approach for sentence-level, word-level, and phrase-level QE that does not require manually extracted features. In 2017, Kim et al. [34] presented a more in-depth study and named it the Predictor-Estimator (PredEst) model, which is a method to address the issues of expensive QE annotations and limited annotated QE data. It consists of two components: the predictor and the estimator. The predictor is a neural word prediction model trained using parallel corpora. It masks the target word, feeds the source language and the corrupted target language into bidirectional RNN (Bi-RNN), and predicts the probability distribution of the masked word. On the other hand, the estimator is a neural QE model trained on QE data, extracting QE feature vectors (QEFVs) and training them on a feedforward network. QEFVs are processed by FNN, RNN, or Bi-RNN to obtain hidden representations, which are then used to predict quality labels for sentence, phrase, or word-level tasks. Later, to train the model effectively, Kim et al. [35] introduced stack propagation and multi-level task algorithms to improve the original method. In 2018, Ive et al. [36] proposed the deepQuest framework for sentence-level and document-level QE, marking the first purely neural approach for document-level QE, which for the first time attempted to experiment with the outputs of both SMT and NMT. Upon testing, the framework proved to be faster and more cost-effective, and greatly improved the performance of the document-level QE framework.
@@
Martins et al. [37, 38] introduced a STACKEDQE system that stacked both linear and neural systems in the WMT17 word-level QE task, followed by combining APE with word-level QE to create an APEQE system. Ultimately, they merged these two systems to form the FULLSTACKEDQE system tailored for word-level QE and extended the FULLSTACKEDQE to sentence-level QE. These systems all achieved commendable results. Building on the approach of Martins et al. [37, 38], Hokamp et al. [39] incorporated features that had been demonstrated to be effective for word-level QE into the input of an NMT system, thus proposing the APE-QE model. This unified APE with word-level QE models and achieved the best performance of the time in both APE and QE tasks.
@@
As the Transformer [40] model has garnered significant success in the field of MT, Fan, Wang, et al. [41] developed a Bilingual Expert model based on a bidirectional Transformer and a PredEst architecture that includes a word prediction module and a QE module. The word prediction module leverages prior knowledge obtained from pretraining on a large parallel corpus and the joint latent representations between the source language and the translation for token prediction, extracting a set of features. Then, they introduce mismatch features that measure the discrepancy between the prior knowledge obtained from well-trained Bilingual Expert and the targets in the QE dataset to train the QE module, which uses a bidirectional LSTM model, achieving SOTA performance at that time. Wang et al. [42] employed a pre-trained Transformer as the predictor and integrated Bottleneck Adapter Layers (BAL) for efficient transfer learning, with specific classifier and regressor as the estimator. They also conducted joint training for word and sentence-level tasks using a unified model and proposed a pseudo PE assisted QE method. This demonstrated the effectiveness of using pre-trained NMT models for transfer learning in QE tasks.
@@
However, Cui et al. [43] argued that the gap between data quality and training objectives in the PredEst framework hindered its ability to benefit from parallel corpora. Consequently, they proposed a framework called DirectQE, which includes a generator for creating pseudo QE data and a detector pre-trained with these pseudo data. This framework allows for pre-training using large parallel corpora and fine-tuning on real QE data, thereby addressing the issues inherent in the PredEst framework.
@@ III-B2 Incorporating Pre-trained Language Models Methods
With the emergence and development of pre-trained LMs such as ELMo [44], BERT [45], XLM [46], and XLM-R [47], some studies have begun to integrate pre-trained LMs into QE models. This integration has enabled better extraction of quality vectors from source texts and translated texts, thereby enhancing the performance of QE systems.
@@
Kepler et al. [48] expanded OpenKiwi [49] into a Transformer-based PredEst model, replacing the predictor with the pre-trained LMs BERT and XLM, and proposed an ensemble method using the POWELL technique to combine word-level and sentence-level predictions. Moreover, they suggested a simple technique for converting word labels into document-level predictions. Wu et al. [50] extended the OpenKiwi by integrating PredEst models based on XLM and Transformer in their submission to WMT20. The former predictor generates both masked and nonmasked representations, while the latter produces only nonmasked representations. The estimator is trained using either LSTM or Transformer, employing top-K and multi-head attention strategies to enhance sentence feature representation. Ranasinghe et al. [51] proposed TransQuest, a PredEst model aimed at reducing the dependence of sentence-level QE on large-scale parallel corpora. TransQuest does not use parallel data to pre-train the predictor but instead directly employs SOTA cross-lingual embedding models like XLM-R [47, 52] to encode the source and target sentences. It consists of two neural networks: MonoTransQuest (MTransQuest) and SiameseTransQuest (STransQuest). MTransQuest uses a single XLM-R model to encode concatenated source and target sentences, whereas STransQuest adopts a Siamese architecture, using separate XLM-R models for the source and translation. Both models use mean squared error loss as the objective function and have shown improved results with specific pooling strategies. Zerva et al. [53] used a pre-trained multilingual encoder combined with adapters, training multilingual models on the OpenKiwi [49] PredEst and found that adapter tuning can resist overfitting. Additionally, they demonstrated that integrating uncertainty information and using out-of-domain data for pre-training can improve QE system performance.
@@
Zhou et al. [54] primarily investigated the application of pre-trained translation models in QE and compared the effectiveness of a bilingual expert, ELMo, and BERT on QE tasks. Yankovskaya et al. [55] contrasted two approaches: one using only BERT and LASER [56] embeddings as features, and the other additionally incorporating log probability features from MT systems. Their research demonstrated the importance of the log probabilities from MT systems.
@@
In 2020, Rei et al. [57] introduced COMET, a neural framework for training multilingual and adaptable MT evaluation models, which is often used for reference-based evaluation to generate predictive estimates of human judgments such as HTER, DA, and MQM. The COMET framework supports two different architectures: an Estimator model and a Translation Ranking model, both consisting of a cross-lingual encoder and pooling layers, with the fundamental difference being the training objective. The Estimator model, which is the most commonly used, is trained to directly regress to a quality score, while the Translation Ranking model is trained to minimize the distance between a “better” hypothesis and its corresponding reference translation and source language.
@@
In 2022, Rei et al. [58] combined the strengths of COMET and OpenKiwi [49] by connecting COMET with the PredEst architecture of OpenKiwi, and equipped it with a word-level sequence tagger and explanation extractor, forming COMETKIWI for QE. COMETKIWI pre-trained the model on metrics data with the learning objective proposed by UniTE model [59], which incorporates reference translation into the training, serving as a form of data augmentation. Additionally, COMETKIWI proposed an interpretability method using attention and gradient information and further refined the influence of attention heads on predictions through a Head Mix module. COMETKIWI also demonstrated the effectiveness of few-shot learning, achieving significant improvements in model performance with only 500 samples.
@@ III-C1 Direct prediction based on content generated by LLMs
Kocmi and Federmann [5] proposed GEMBA, a GPT-based [60] translation quality assessment metric that uses single-step prompting and can be applied to scenarios with reference translations as well as to QE. They evaluated 9 different GPT models and concluded that only GPT-3.5 and larger models are capable of performing QE. GEMBA focuses on zero-shot prompting, with the authors employing 4 different prompt templates to execute quality assessments for both reference-based and nonreference-based translation modes. GEMBA predicts scores directly based on content generated by LLMs, assessing each segment independently, then averaging the scores of all segments to obtain a final system-level score, achieving SOTA performance at the system level, but lacking in segment-level analysis. To enhance the performance of LLMs in quality assessment, Lu et al. [6] introduced Error Analysis Prompting (EAPrompt), a new prompting method that combines Chain-of-Thought (CoT) [61] with EA [62]. Using ChatGPT, this approach predicts the degree and number of errors and provides a score based on the severity of those errors, creating MQM-like assessments. It achieved better results on GPT-3.5-turbo than GEMBA. Yang et al. [7] introduced the Knowledge-Prompted Estimator (KPE), which is a CoT approach that integrates three single-step prompting techniques, using LLMs to predict fluency, word-level similarity, and sentence-level similarity [63], resulting in better performance for segment-level QE. Furthermore, KPE also demonstrated its advantages in terms of interpretability.
@@ III-C2 Based on the generative probabilities of LLMs
Huang et al. [8] proposed an unsupervised QE framework that can explore useful information in LLMs. This framework makes use of both the sequence-level probability and the uncertainty of LLMs as evidence for how well LLMs perform QE. Specifically, they designed translation style prompts to guide LLMs in generating sequence-level probability, while they also investigated the impact of LLM’s uncertainty on sequence-level probability. This was done by designing various prompts and demonstrations, and feeding the same sample into the LLMs multiple times in different forms, which resulted in multiple distinct generation probabilities. Through their analysis, they found that prompts of translation style and successful demonstrations can enhance the performance of the model.
@@ III-C3 Leveraging LLMs to generate pseudo data
Xu et al. [9] introduced INSTRUCTSCORE, a method that learns interpretable text generation metrics without the need for human-annotated scores. This method constructs MQM-like data using knowledge provided by GPT-4 [64] to train the LLaMA model [65]. First, they use GPT-4 to build pseudo data with errors and explanations, which is then used for fine-tuning LLaMA model. After that, they sample and input real-world diagnostic output into GPT-4 for feedback, and then fine-tune the LLaMA model further by choosing the explanation most consistent with human input. They repeat these latter two steps to optimize the model’s output. Ultimately, they found that the model performs well across tasks, domains, dimensions, and even on unseen tasks. Huang et al. [10], on the other hand, leveraged LLMs to corrupt reference sentences. They then generated a fluent sentence from the corrupted one and used this as a noisy negative view. This approach, since it does not require data annotation, has strong generalization capabilities.
@@ III-C4 LLMs as the foundation for QE models
Gladkoff et al. [11] fine-tuned LLMs using the OpenAI API interface to assess whether translations require post-editing. Their results showed that increasing the size of LLMs does not significantly improve the performance of the evaluation task. Similarly, Xu et al. [9], as mentioned earlier, also utilized pseudo data generated by GPT-4 [64] to train the LLaMA model [65].